### v1

# GINTRIP: Interpretable Temporal Graph Regression using Information bottleneck and Prototype-based method

Abstract
Deep neural networks (DNNs) have demonstrated remarkable performance across various domains, yet their application to temporal graph regression tasks faces significant challenges regarding interpretability. This critical issue, rooted in the inherent complexity of both DNNs and underlying spatio-temporal patterns in the graph, calls for innovative solutions. While interpretability concerns in Graph Neural Networks (GNNs) mirror those of DNNs, to the best of our knowledge, no notable work has addressed the interpretability of temporal GNNs using a combination of Information Bottleneck (IB) principles and prototype-based methods. Our research introduces a novel approach that uniquely integrates these techniques to enhance the interpretability of temporal graph regression models. The key contributions of our work are threefold:
We introduce the Graph INterpretability in Temporal Regression task using Information bottleneck and Prototype (GINTRIP) framework, the first combined application of IB and prototype-based methods for interpretable temporal graph tasks. We derive a novel theoretical bound on mutual information (MI), extending the applicability of IB principles to graph regression tasks. We incorporate an unsupervised auxiliary classification head, fostering multi-task learning and diverse concept representation, which enhances the model bottleneck’s interpretability.
Our model is evaluated on real-world traffic datasets, outperforming existing methods in both forecasting accuracy and interpretability-related metrics. The implementation of our research is available at Github.
Index Terms:
Temporal Graph, Regression, Interpretability, Information Bottleneck, Prototype
00footnotetext: Equal contribution
Figure 1: Proposed method architecture. It includes an encoder, learnable prototypical layer, subgraph extractor, regression, and classification head components.
I Introduction
In the intelligent transportation industry, ensemble tree-based methods such as XGBoost [1] are in high demand. Their popularity is due to their interpretability. Meanwhile, with the success of Graph Neural Networks (GNNs) in a wide range of deep learning tasks, there has been an increasing demand to explore the decision-making process of these models and provide explanations for their predictions. Existing traffic dashboards such as STRADA[2] cannot interpret the predictions. To meet this requirement, explainable AI (XAI) has developed as a means to interpret black-box models by offering clear explanations for their outputs.
Interpretability in deep neural networks (DNNs) is one of the most important neglected aspects that still need more investigation due to the black-box nature of DNNs. Different tools and methods have been used to address this black box. One of them is the IB [3] that is raised when analyzing DNNs using information theoretic concepts. On the other hand, some other techniques, like prototype-based methods, use learnable blocks that can capture key features that are useful for interpretability[4, 5]. For instance, [6] attempted to optimize IB bounds via adversarial training for semantic segmentation tasks while the interpretability is yet neglected. On the other hand, [7] tackles the same task with the notion of prototypes, leading to explainability.
In the area of graphs, there are prominent works to integrate IB. Specifically, [8, 9] are the pioneers who derive tractable bounds to achieve interpretable Graph Neural Networks (GNNs). In the terminology of interpretability, they are classified into the post-hoc methods, which help to understand and interpret the predictions of black-box models like deep neural networks [10],[11], which are often opaque in terms of decision-making processes. On the other hand, another class of interpretable models are self-explainable methods. One such approach is to utilize the IB principle to explore the natural interpretability and generalization capabilities of GNNs [12]. Furthermore, recent progress in intrinsic approaches [13] has tackled explainability issues and addressed challenges in managing graph out-of-distribution cases by applying invariant learning and causal inference techniques. In the temporal graph domain, IB methods are unexplored. One of the main works in this area is [14], which generates spatial and temporal subgraphs as explainability output. These subgraphs are not linked to prototypes, hindering high-level interpretability.
In this study, we tackle the aforementioned challenges by presenting the GINTRIP framework. Our framework offers temporal graph prediction and interpretability as a self-explainable model. To the best of our knowledge, we are the first to propose this setting. In summary, our contributions are as follows:
  • We are the first model to perform interpretable temporal graph regression based on the prototype notion
  • we provide pseudo-labels to facilitate interpretability
  • Our framework derives tractable MI bounds for the temporal graph regression problem.
  • Our model is extensively evaluated on real-world traffic datasets, for which it outperforms the state-of-the-art.
II Methodology
II-A Problem Statement
We focus on temporal graph regression, a technique for predicting time-dependent node values in graph-structured data. We apply this to traffic forecasting, which anticipates traffic conditions based on historical data. In our graph model $G_{in}=(V_{G_{in}},E,A)$. Here, $V_{G_{in}}=\left\{V_{1},V_{2},...,V_{N}\right\}$ is the node set, $E$ the edge set, and $A\in\mathbb{R}^{N\times N}$ the adjacency matrix. At each timestamp $t$, the graph $G_{in}$ includes a dynamic feature matrix $X_{t}\in\mathbb{R}^{N\times D}$. In this study, the graph is static and the traffic feature is speed. Given a length-${T^{\prime}}$ forecast horizon and a length-$T$ observed series history $\left[y_{1},...,y_{T}\right]\in\mathbb{R}^{T}$, the goal is to predict the vector of future values. For simplicity, we consider a sliding window of length $W\leq T$ ending with the most recent observed value $y_{T}$ as the model input, denoted as $\boldsymbol{x}\in\mathbb{R}^{W}=\left[y_{T-W+1},y_{T-W+2},...,y_{T}\right]$. The prototype layer is also incorporated into our model. It is defined as a set ${G}_{p}=\left\{\mathbf{z}_{G_{p}}^{1},\mathbf{z}_{G_{p}}^{2},\cdots,\mathbf{z}_{G_{p}}^{M}\right\}$, where $M$ is the total number of prototypes, and each prototype $\mathbf{z}_{G_{p}}^{m}$ is a learnable parameter vector that acts as the latent representation of the prototypical part (i.e., ${G}_{p}$) of graph ${G_{in}}$. We allocate $J$ prototypes for each pseudo-class, i.e., $M=K\times J$, where $K$ denotes the number of classes corresponding to our unsupervised pseudo-labels. Indeed, for each node $V_{i}$, a threshold $T_{i}=Q_{0.1}(X_{:,i,:})$ classifies each node into one of our predefined pseudo-classes(congested/non-congested).
II-B Graph information bottleneck
The MI between two random variables of $X,Y$ is defined as follows:
$\small I(X;Y)=\int_{X}\int_{Y}p(x,y)\log\frac{p(x,y)}{p(x)p(y)}\,dx\,dy$ (1)
The IB method creates a compressed representation $Z$ from input $X$, preserving essential components for predicting output $Y$ while discarding redundant information. The IB objective is:
$\small\min_{Z}-I(Y;Z)+\beta I(X;Z)$ (2)
Where $\beta$ is the Lagrange multiplier that controls the trade-off between two terms. The first part is responsible for the compression concept, and the second concerns output prediction.
The IB concept has recently been found to be applicable in graph learning, resulting in the development of the IB-Graph. This approach aims to create a condensed representation of a given graph $G_{in}$ while preserving its essential characteristics. The method draws inspiration from the Graph Information Bottleneck (GIB) principle, which focuses on extracting a compact yet interpretable subgraph $G_{sub}$ from the original graph $G$. This extraction process is guided by an optimization objective designed to balance information retention and compression:
$\small\min_{G_{sub}}-I(Y;G_{sub})+\beta I(G_{in};G_{sub})$ (3)
II-C Prototype Injection
Prototype-based methods are used in different machine-learning tasks to capture the essential key of the bottleneck variable of the input. Combining this technique with GIB could provide and extract a more meaningful subgraph of the input graph. Ultimately, we develop the prototype layer as a learnable component to discern predictive temporal patterns. To do so, we compute the similarity of the extracted subgraph and the prototype matrix as a predictive feature during forecasting. Using $G_{p}$, we can inject $G_{p}$ trainable variables to the first term of (3) i.e. $I(Y;G_{sub})$. Based on the chain rule of MI, we can write: $I(Y;G_{sub})=I(Y;G_{sub},G_{p})-I(Y;G_{p}|G_{sub})$ , and then we can rewrite (3): $\small\min_{G_{sub}}-I(Y;G_{sub},G_{p})+I(Y;G_{p}|G_{sub})+\beta I(G_{in};G_{sub})$. Due to some difficulty in computing the second term, we again use the chain rule of MI. We can decompose the second term, and we have a more extended expression:
$\min_{G_{sub}}$ $-I(Y;G_{sub},G_{p})+I(G_{p};Y,G_{sub})-I(G_{p};G_{sub})$ (4)
$+\beta I(G_{in};G_{sub})$
Observing the first two terms in (4), we obviously recognize that they are opposite, meaning the first MI wants to increase, and the second one wants to decrease. But, after enough training time, these two terms force each other, where an increase in $I(Y;G_{sub},G_{p})$ leads to an increase in the $I(G_{p};Y,G_{sub})$ that is different behavior than (4). So, we exclude the second MI term from our optimization objective to prevent this misalignment:
$\small\min_{G_{sub}}-I(Y;G_{sub},G_{p})-I(G_{p};G_{sub})+\beta I(G_{in};G_{sub})$ (5)
After sufficient training, $I(G_{p};G_{sub})$ increases, indicating stronger correlation between $G_{p}$ and $G_{sub}$. Fig. 1 shows the relationship complexity of $G_{p}$ and $G_{sub}$ is lower than that of $G_{sub}$ and $Y=(Y_{cls},Y_{reg})$. Assuming $G_{p}\approx G_{sub}$, we get $I(G_{p};G_{sub})\approx H(G_{sub})$, leading to $I(G_{p};G_{sub})\gtrsim I(Y;G_{sub})$. Consequently, $I(G_{p};Y,G_{sub})\gtrsim I(Y;G_{sub},G_{p})$, implying an increase in $I(Y;G_{sub},G_{p})$ leads to an undesirable increase in $I(G_{p};Y,G_{sub})$, conflicting with optimization (4).
II-D Subgraph Extraction
Our model adopted Spatio-Temporal Graph Neural Networks (STGNN) to process the input temporal graph. More specifically, we used MTGNN[15] as our GNN backbone. Following [16], we passed the nodes’ extracted features, referred to as $h_{i}$ in Fig. 1, to the sequence of an MLP layer and a sigmoid activation function. Their output assigns the probability $p_{i}$ of the node inclusion in the selected subgraph, output of the $G_{sub}$-Extractor block in Fig. 1. The final representation of each selected subgraph node would be: $\small z_{i}=\lambda_{i}h_{i}+(1-\lambda_{i})\epsilon$, where $\lambda_{i}\sim Bernoulli(p_{i})$ and $\epsilon\sim\mathcal{N}(\mu_{h_{i}},\sigma_{h_{i}}^{2})$. The learned probability $p_{i}$ facilitates selective information retention in $G_{sub}$. This approach preserves interpretability within the subgraph while also potentially aiding the learning of prototypes introduced in the previous step. To minimize the MI between the input graph and the selected subgraph, [17] showed that it is equivalent to minimizing the following upper bound for $I(G_{in},G_{sub})$:
$\mathbb{E}_{G_{in}}\left[-\frac{1}{2}\log S+\frac{1}{2|V_{G_{in}}|}S+\frac{1}{2|V_{G_{in}}|}M^{2}\right]\triangleq\mathcal{L}_{sub}$ (6)
where $S=\sum_{i=1}^{|V_{G_{in}}|}(1-\lambda_{i})^{2}$ and $M=\frac{\sum_{i=1}^{|V_{G_{in}}|}\lambda_{i}(h_{i}-\mu_{h_{i}})}{\sigma_{h_{i}}}$. After noise adding, we compute the embedding $Z_{sub}$ by adopting mean pooling on the subgraph[18].
To enforce learning connected subgraphs[16], we employ the following loss function, which is called connectivity loss: $\small\mathcal{L}_{con}\triangleq\left\|\widehat{P^{T}AP}-I_{2}\right\|_{F}$. where $P\in\mathcal{R}^{\left|V_{G}\right|\times 2}$ and $A$ are the node probability assignment and adjacency matrix at the batch level, respectively. $I_{2}$ is the 2x2 identity matrix, $\|...\|_{F}$ is the Frobenius norm operation and $\widehat{(...)}$ is the row normalization. More precisely, each $P$ matrix row is related to the probability of its corresponding node, meaning that for each row we have ($p_{i},1-p_{i}$). Notably, locality preservation is a crucial aspect of the extracted subgraph[19].
[…]
II-E Theoretical Bounds
To optimize our objective, we must consider the inherent characteristics of each MI term. By applying appropriate lower and upper bounds to these terms, we define a constrained optimization problem that forms our final loss function. Below, we derive lower bounds for the first two terms of the above optimization objective. $I(Y_{cls};G_{sub},G_{p})$ is equal:
$\mathbb{E}_{Y_{cls},G_{sub},G_{p}}[\log p(Y_{cls}|G_{sub},G_{p})]-\mathbb{E}_{Y_{cls}}[\log p(Y_{cls})]$ (7)
$\geq\mathbb{E}_{Y_{cls},G_{sub},G_{p}}[\log p(Y_{cls}|\gamma(G_{sub},G_{p}))]-\mathbb{E}_{Y_{cls}}[\log p(Y_{cls})]$
$\geq\mathbb{E}_{Y_{cls},G_{sub},G_{p}}[\log q_{\phi}(Y_{cls}|\gamma(G_{sub},G_{p}))]\triangleq-\mathcal{L}_{cls}$
Where $\gamma$ is a similarity function between $G_{sub},G_{p}$. The first inequality was derived from DPI (Data Process Inequality)[20] and the second one, caused by $q_{\phi}(Y_{cls}|\gamma(G_{sub},G_{p}))$, which is the variational approximation to the true posterior $p(Y_{cls}|\gamma(G_{sub},G_{p}))$ Equation (7) illustrates that maximizing the MI $I(Y_{cls};G_{sub},G_{p})$ is equivalent to minimizing the classification loss, $\mathcal{L}_{cls}$. In practice, the cross-entropy loss is typically employed for categorical $Y_{cls}$.
The previous derivation can be extended to our regression head, requiring additional mathematical (using Pinsker’s inequality) steps. This extension results in a lower bound expressed in terms of MSE:
$I(Y_{reg};G_{sub},G_{p})\geq-\lambda MSE(\hat{Y}_{reg},Y_{reg})\triangleq-\lambda\mathcal{L}_{reg}$ (8)
Where $\lambda$ is a constant coefficient and also $\hat{Y}_{reg}$ and $Y_{reg}$ are the classification head prediction and the classification pseudo-label (ground-truth), respectively (refer to Fig. 1).
Following a similar variational approach to that used in equation (8), we can express:
$I(G_{sub};G_{p})\geq-\lambda^{\prime}MSE(Z_{sub},Z_{G_{p}})\triangleq-\lambda^{\prime}\mathcal{L}_{var}$ (9)
Here, $\lambda^{\prime}$ is a constant coefficient. $Z_{sub}$ denotes the output of the pooling function applied to $G_{sub}$, while $Z_{G_{p}}$ represents the concatenation of all prototype vectors $\left\{\mathbf{z}_{{G}_{p}}^{1},\mathbf{z}_{{G}_{p}}^{2},\cdots,\mathbf{z}_{{G}_{p}}^{M}\right\}$. For $q_{\phi}$, we employ a single-layer perceptron to estimate $Z_{G_{p}}$ from $Z_{sub}$. The total loss function is a convex combination of individual losses derived from the aforementioned bounds: $\mathcal{L}=\lambda_{1}\mathcal{L}_{reg}+\lambda_{2}\mathcal{L}_{sub}+\lambda_{3}\mathcal{L}_{var}+\lambda_{4}\mathcal{L}_{con}+\lambda_{5}\mathcal{L}_{cls}$. To address the instability of this function during training, we employed the multi-loss variation coefficient adjustment approach [21].
III Experiments
We evaluate the performance of GINTRIP in terms of forecasting metrics and explainability on three distinct real-world traffic datasets. In terms of explainability metrics, the fidelity and sparsity are defined as[16]:
$\small\text{Fidelity}_{+/-}(\%)=\frac{100}{\lfloor T/W\rfloor}\sum_{i=1}^{\lfloor T/W\rfloor}\left(|f(G_{i})-f(G_{i}^{+/-})|\right)$
$f$ represents the trained predictive spatio-temporal function. $G_{i}^{-}$ and $G_{i}^{+}$ refer to the masked spatial-temporal complementary subgraph and the selected subgraph structure for $i^{th}$ $G_{in}$, respectively. It directly measures the quality of the learned prototypes on the targeted task. $k$ represent the number of important nodes in the original graph, denoted as Sparsity. Node importance is determined based on the assignment probability explained in Section II-D.
III-A Quantification analysis
We conducted our experiments on real-world traffic datasets. Specifically, PeMS04, PeMS07, and PeMS08 [22, 23] were chosen as graph-based traffic datasets.
To evaluate the proposed method, we compare GINTRIP with two groups of methods:
  1. 1.
    STGNN Traffic Methods:
    • STSGCN[24], DSTAGNN[25]: employs gated spatial-temporal graph convolution to proceed the traffic as a spatio-temporal tensor by different aggregation techniques.
  2. 2.
    Explainable STGNNs: STExplainer[14]: decomposes the explainability into the spatial and the temporal domain.
In Table I, the traffic forecasting metrics are presented. The results indicate that self-explainablity promotes the predictivity capacity. In contrast to [14], our extracted subgraphs are time-aware, which eases the interpretation process. Indeed, it dispenses with the post-processing of temporal and spatial subgraphs proposed in [14].
TABLE I: Traffic predictions benchmark, B.B. is the abbreviation for block box models which are not interpretable
Baselines PeMS04 PeMS07 PeMS08
[…]
B.B. STGNCDE [24] 19.21 31.09 12.76 20.53 33.84 8.80 15.45 24.81 9.92
DSTAGNN [25] 19.30 31.46 12.70 21.42 34.51 9.01 15.67 24.77 9.94
Interpretable STExplainer [14] 19.09 30.64 12.21 20.00 33.45 8.51 14.70 23.91 9.80
GINTRIP (Ours) 18.62 29.02 12.75 20.12 33.53 8.62 14.88 24.05 9.89
To verify our interpretability, we compare the fidelity and sparsity with STExplainer[14], which is presented in Fig.2. Our fidelity score($F_{+}$) is superior to the previous state-of-the-arts.
Figure 2: Fidelity over Sparsity. High fidelity is an indicator of interpretability
III-B Visualizations
Fig. 3 illustrates the interpretability of the learned model by presenting extracted subgraphs at various sparsity thresholds ($k$). These subgraphs, derived from the Pems04 dataset, exhibit consistent spatio-temporal patterns across the selected subgraph nodes regarding different $k$ values. This consistency underscores GINTRIP’s ability to extract robust, time-aware subgraphs using the IB technique and learned prototypes.
Figure 3: Positive over the negative fidelity based on the two selected subgraphs regarding the sparsity (k)
IV Conclusion
In this study, we emphasize the significance of explainability in spatio-temporal graph neural networks. To address this, we propose a novel framework called GINTRIP that not only predicts future spatio-temporal signals accurately but also provides interpretable time-aware subgraphs. Our framework incorporates the prototype learning approach which employs variational approximation for tractability. Additionally, we introduce a unified Temporal-GNN that provides intrinsically explainable and robust spatio-temporal representations. Through comprehensive experiments, we examine the hypothesis that self-explainability causes predictability and robustness. Our results surpass existing state-of-the-art baselines in both predictive accuracy and explainability. In future research, we plan to investigate interpretability in hypergraphs.
Acknowledgment
This work is funded by Innoviris within the research projects TORRES. AR and LDC acknowledge support from the Belgian Science Policy Office (BELSPO) under contract number B2/233/P2/PRECIP-PREDICT and through the FED-tWIN programme (Prf-2020-017).